question 2

hello can you help me with these Q1 AND Q2

THAN YOU

2. An account in Building Society A pays 7.4% interest, compounded quarterly. while another Building Society B pays 7.5% with annual compounding. You have £1,000 to invest. What is the value of your investment in these Societies after three years? Which Society do you prefer? What additional amount would you have to invest in this Society at the start of the fourth year to achieve a future value of £2,000 at the end of the fourth year?
9. The Board of Directors of FDM Pic are considering the following projects (all values in Es) Year Project 2 200,000 Project 3 -100,000 30,000 Project Project 4 0 -150,000 -100,000 60,000 10,000 2 80,000 50,000 30.000 40,000 50,000 4 90,000 10,000 340,000 Given the riskiness of these projects, investors require a 15% return on projects 1 and 2 and a 20% return on projects 3 and 4 a) Calculate the payback period for each project and state what decision FDM will reach if they use a three-year payback period (7 marks) b) Calculate the IRRs for projects 1 and 4. What is the appropriate accept/reject decision for these two projects? (7 marks) c) Calculate the NPV of each project. Which projects should FDM accept? (7 marks) d) Discuss the implications of your answers in parts (a) and (c) (7 marks) e) Discuss the implications of your results in parts (b) and (c) if projects1 and 4 are mutually exclusive (7 marks)
Answer #1

I will help you out with the first question. The second question is a part of graded homework / exam / quiz. As per honor code, we will not be able to help you out with that.

First question:

FV = PV x (1 + i)n where i = interest rate per period and n = number of periods

Society A,

One period is one quarter.

Hence, n = nos. of quarters in 3 years = 4 x 3 = 12

i = 7.4%/4 = 1.85%

Investment value after 3 years = FV = 1000 x (1 + 1.85%)12 =  1,246.04

Society B,

One period is one year. Hence, n = 3 and i = 7.5%

Hence, investment value after 3 years, FV = 1000 x (1 + 7.5%)3 =  1,242.30

FV is marginally better in case of Society A. Hence, I will prefer Society A.

Let's assume the incremental investment at the beginning of year four be P. Then,

(1,246.04 + P) x (1 + 1.85%)4 = 2,000

Hence, P = 2,000 / 1.01854 - 1,246.04 =  612.56

Hence, the incremental investment required at the beginning of year four = P = 612.56